# Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images

### v2

@@ 1 Introduction
Producing compelling 3D assets has become an increasingly active area of research in the field of generative AI. Despite the significant progress in training large-scale generative models of 3D geometry [LGT23, LWVH23, WLW24, GSW22, JN23, NJD22, VWG22, LXJ24, LSC24, SCZ23, ZWZ24], synthesizing compelling, seamless, and high-quality textures for 3D shapes and scenes still remains challenging. One major obstacle is the limited availability of training 3D asset datasets with high-quality textures. Several recent approaches [CSL23, RMA23, CKF23, CDG24, ZCQ24, LXLW24, CMZ24] have resorted to leveraging powerful generative 2D image models, such as denoising diffusion models pre-trained on massive 2D datasets, to guide surface texture generation. These approaches generate images across different views conditioned on text prompts and depth maps rendered from the given 3D shapes or scenes, then attempt to combine the generated images onto the surface.
@@
Unfortunately, these methods suffer from a number of limitations. First, merely projecting the generated images back to the surface tends to generate texture distortions in areas of rapid depth changes or high curvature surface regions (Figure Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images, left). Second, the images generated from different views are often combined with ad-hoc criteria to create a single surface texture. For example, many methods [CSL23, RMA23] simply transfer the colors from the most front-facing view to each texel (i.e., pixel in the texture map), causing visible seams in the texture maps especially in areas where colors of neighboring texels are copied from different views (Figure Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images, middle & right). Other methods rely on simple averaging schemes of RGB colors [CDG24] or image latents, causing significant color bleeding. Others resort to global texture optimization techniques [CKF23, LXLW24], which are slow and can still fail to generate coherent textures since their initialization is still based on ad-hoc thresholds for view selection and color averaging.
@@
Our approach, named  Im2SurfTex, addresses the above limitations with the introduction of a novel, optimization-free module that can be easily integrated to existing texture generation approaches. The module is trained to combine color information from multiple viewpoints to textures through a cross-attention mechanism, where for each surface point, several candidate image neighborhoods across different views are examined and combined back in texture space depending on local surface geometry, as encoded in 3D positions, normals and geodesic distances within these patches. In this manner, the attention mechanism captures local context based on surface (geodesic) proximity rather than relying solely on 3D Euclidean proximity that might correlate surface region textures far from each other in geodesic sense. Our module yields coherent textures efficiently, without requiring any slow optimization procedures. Our experiments indicate significant improvements in generated texture quality, measured by different scores, including FID [HRU17], KID [BSAG18], CLIP [SBV22] metrics, when compared to alternatives.
@@ 2 Related Work
Classic texture generation methods mostly focused on example-based approaches [WLKT09], including region-growing techniques [EL99, WL00] and strategies leveraging local coherence [Ash01, TZL02] to maintain consistency across synthesized textures. Patch-based methods [EF01, KSE03] synthesized textures using patch patterns from reference images, while other techniques [KEBK05, HZW06] progressively refined synthesized texture based on optimization procedures. Another significant research direction involved directly generating textures on 3D surfaces [Tur01, ZMT06, FSDH07] by exploiting vector fields defined over the surface to seamlessly map textures onto complex geometries. All these example-based approaches were unable to capture texture variability, generate diverse textures, or handle diverse shapes.
@@
Our method builds upon diffusion models [SDWMG15, HJA20, ZZZ23], which have demonstrated superior performance compared to GANs [GPAM14, ZPIE17] in image generation tasks [RBL22, NDR21, SCS22, RDN22, ZRA23]. Closely related to our approach are text-guided generation models for 3D object synthesis, where text-to-image diffusion models are used for distilling 3D objects as neural radiance fields [MST20, KKLD23] via Score Distillation Sampling [PJBM22, WDL23]. Following DreamFusion [PJBM22], several approaches have been proposed [LGT23, MRP23, CWWL24, WLW24, TMT24, SWY24]. However, these methods do not specifically target the task of texture generation.
@@
Initial efforts in texture generation via text-to-image diffusion models, such as Text2Tex [CSL23] and TEXTure [RMA23], employed depth-conditioned diffusion models [RBL22, ZRA23] to iteratively inpaint and refine the textures of 3D objects. Both methods start with a preset viewpoint, generating texture updates for corresponding regions of the 3D object by back-projecting depth-guided views. In Text2Tex, a coarse texture is progressively created by iterating over multiple viewpoints and refining the texture map based on high- surface-coverage viewpoints. This refinement applies a denoising diffusion process of moderate strength to preserve the texture’s original appearance while enhancing details. Similarly, TEXTure divides the texture map into distinct regions labeled as keep, refine, or generate, enabling selective refinement or generation of textures. Despite these efforts to achieve global consistency, these methods employ ad hoc thresholds to define the different shape regions and hand-engineered strategies for backprojection, often leading to seams between texture regions synthesized from different viewpoints.
@@
To address these issues, other methods such as TexFusion [CKF23] leverage latent diffusion to interlace diffusion and back-projection steps, producing 3D-aware latent images that are subsequently decoded and merged into a texture map. Similarly, SyncMVD [LXLW24] employs a latent texture map where all views are encoded at each denoising step, further enhancing consistency in geometry and appearance. TexGen [HGZ24] introduced a multi-view sampling and resampling framework that updates a UV texture map iteratively during denoising, aiming to reduce view discrepancies. Still, these methods rely on ad hoc blending masks or heuristics for aggregating texture information from different views, such as mere averaging or using the most front-facing view information for each texel. In contrast, our approach learns to aggregate color information from multiple views based on both geometry and texture information, promoting the generation of more coherent and seamless surface texture maps.
@@
A more recent approach, Paint3D [ZCQ24], achieves impressive texture generation results, by adopting a two-stage texture generation strategy. In the first stage, a coarse texture is created by backprojecting views to texture space via the heuristic of using the most front-facing view information for each texel, as in previous methods. The second stage involves a refinement and inpainting process that utilizes a diffusion model in UV texture space, conditioned on a UV position map encoding the 3D adjacency information of texels. While this method directly encodes 3D geometric information into the texture map, it can still result in misaligned textures due to its employed heuristic during its coarse stage. The subsequent texture refinement steps often fail to fix the artifacts of the coarse stage, as demonstrated in our experiments. Another related method, MatAtlas [CDG24], incorporates a three-step denoising process with sequential operations and line conditions to preserve geometry and style consistency. However, MatAtlas employs an averaging heuristic for blending the final texture from the generated views, leading to inconsistencies or overly smooth surfaces, as also demonstrated in our experiments. TEXGen [YYG24] takes a different approach by directly training a large-scale diffusion model in the UV texture space, and integrating convolution operations in UV space with 3D-aware attention layers in their denoising network to achieve high-resolution texture synthesis. However, their method still faces challenges in maintaining cross-view consistency since the generated textures are conditioned on single-view images which are merely backprojected to the UV space to derive the initial partial texture maps used in their diffusion.
@@
In a concurrent work, MVPaint [CMZ24] introduces a multi-stage texture generation framework. In the initial stage, a latent texture map is employed during multi-view projection to create a synchronized texture across multiple views, similar to SyncMVD [LXLW24]. This is followed by an inpainting stage, where uncovered texture regions are filled using a dense colored point cloud extracted from the generated texture map. Colors are propagated to empty texels in a spatially aware manner using inverse distance weighting and normal similarity between neighboring points. Finally, a refinement stage upscales the texture map and smooths out seams through weighted color averaging among k-nearest neighbors in 3D space. Still, MVPaint relies on an averaging scheme to aggregate view information into texture space. In contract, our approach learns this aggregation by encoding both geometric and appearance information from multiple views to produce textures with greater consistency.
@@
Expanding beyond single object texture generation, InstanceTex [YGC24] focuses on texture generation for 3D scenes, employing a local synchronized multi-view diffusion strategy to improve local texture consistency across multiple objects. 3D Paintbrush [DLAH24] specializes in localized stylization of single objects, using cascaded score distillation to refine textures within specific object regions. These approaches differ from our method in scope: InstanceTex is tailored for stylistic consistency in large environments, while 3D Paintbrush targets localized edits.
@@
Figure 2: (Top) The Im2SurfTex, pipeline utilizes depth images and a text prompt to generate a number of candidate views (RGB images) for a given shape. The views are aggregated through a learned backprojection module that incorporates geometric information, such as 3D location, normals, angles between normals, and view vectors, as well as geodesic neighborhood information (bottom right) of shape points corresponding to pixels of the generated RGB images. The backprojection module integrates several cross-attention blocks (bottom left) used to infer texel features and colors from the appearance and geometric information gathered from relevant, non-background pixels across all available views. As some texels may remain uncolored, an inpainting and high-definition (HD) module is applied to refine the texture map following Paint3D [ZCQ24].
@@ 3.1 Depth/edge map rendering & viewpoint selection
As done in several recent texture generation approaches [ZCQ24, LXLW24, CDG24, CMZ24], the first step in our pipeline is to render the mesh into a set of depth maps $\{\mathbf{D}_{c}\}_{c=1}^{C}$ from various viewpoints, where $C$ is the total number of viewpoints. These maps are used as conditioning to guide the diffusion process to generate images consistent with the depth cues. There have been various strategies for viewpoint selection and diffusion model conditioning – in our paper, we experimented with two backbones: one based on Paint3D [ZCQ24], and another based on MatAtlas [CDG24], briefly described below.
@@
Paint3D follows an iterative strategy of viewpoint selection, image generation, and backprojection of the generated images to texture space. First, a couple of $1024\times 512$ depth images are generated from the frontal and rear views of the shape and are concatenated in a $2\times 1$ grid. The grid is passed as input to a diffusion process that generates a corresponding $2\times 1$ grid of RGB images. The use of both views as input to the diffusion model helps with the view consistency [ZCQ24]. The generated images are backprojected to the shape texture through a simple inverse UV mapping strategy – we discuss backprojection strategies, including ours, in Section 3.3. The next iteration proceeds with two side-wise viewpoints, from which both depth images and partially colored RGB images are rendered from the partially textured mesh. These are provided as a grid to another diffusion process, whose generated images are again backprojected to UV space. The process repeats for one more step where two other top- and bottom-wise viewpoints are used. In total, three iterations (total $C=6$ views), with two viewpoints processed at a time, yielded the best results in Paint3D. Our experiments with this backbone follow the same iterative procedure and viewpoints – we only modify the backprojection.
@@
MatAtlas [CDG24] follows a different viewpoint selection, diffusion conditioning, and view generation strategy. Initially, a set $400\times 400$ depth maps are rendered from viewpoints uniformly sampled from the viewing sphere, and are arranged in a $4\times 4$ grid. In addition, $16$ edge maps are created using the shape’s occluding and suggestive contours and are placed also in a grid. These two grids are used as input to the diffusion process that generates a $4\times 4$ grid of RGB images. These are backprojected and blended into the shape’s texture space (discussed in Section 3.3). The resulting partially textured shape is rendered from the same viewpoints, and the rendered RGB images along with added partial noise, are passed to a second diffusion process yielding an updated set of RGB views. These are backprojected to the shape’s texture space, yielding a sharper texture [CDG24]. In a third step, additional viewpoints are selected accessing shape regions not textured yet. The textured shape is rendered from these viewpoints and the rendered images are arranged in a grid processed through another diffusion process, which generates another set of RGB images. These are again backprojected to the final shape’s texture. In our implementation, we use $6$ initial viewpoints to render depth maps at resolution $512\times 512$, arranged in a $3\times 2$ grid (we do not make use of edge maps). The used viewpoints are the same as the ones used in Paint3D for more fair comparisons across the two backbones, and also because we observed that texture details are better preserved from the higher resolution depth maps. We replace the MatAtlas backprojection with ours – the rest of the pipeline follows MatAtlas.

### v3

@@ 1 Introduction
Producing compelling 3D assets has become an increasingly active area of research in the field of generative AI. Despite the significant progress in training large-scale generative models of 3D geometry [Lin:2023:Magic3D, Liu:2023:Zero, Wang:2024:Prolificdreamer, Gao:2022:Get3d, Jun:2023:Shape, Νichol:2022:Pointe, Vahdat:2022:Lion, Liu:2024:One, Liu:2023:One2345plus, Shi:2023:Zero123plus, Zhang:2024:CLAY], synthesizing compelling, seamless, and high-quality textures for 3D shapes and scenes still remains challenging. One major obstacle is the limited availability of training 3D asset datasets with high-quality textures. Several recent approaches [Chen:2023:Text2tex, Richardson:2023:Texture, Cao:2023:Texfusion, Ceylan:2024:Matatlas, Zeng:2024:Paint3D, Liu:2024:SyncMVD, Cheng:2024:MVPaint] have resorted to leveraging powerful generative 2D image models, such as denoising diffusion models pre-trained on massive 2D datasets, to guide surface texture generation. These approaches generate images across different views conditioned on text prompts and depth maps rendered from the given 3D shapes or scenes, then attempt to combine the generated images onto the surface.
@@
Unfortunately, these methods suffer from a number of limitations. First, merely projecting the generated images back to the surface tends to generate texture distortions in areas of rapid depth changes or high curvature surface regions (Figure Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images, left). Second, the images generated from different views are often combined with ad-hoc criteria to create a single surface texture. For example, many methods [Chen:2023:Text2tex, Richardson:2023:Texture] simply transfer the colors from the most front-facing view to each texel (i.e., pixel in the texture map), causing visible seams in the texture maps especially in areas where colors of neighboring texels are copied from different views (Figure Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images, middle & right). Other methods rely on simple averaging schemes of RGB colors [Ceylan:2024:Matatlas] or image latents, causing significant color bleeding. Others resort to global texture optimization techniques [Cao:2023:Texfusion, Liu:2024:SyncMVD], which are slow and can still fail to generate coherent textures since their initialization is still based on ad-hoc thresholds for view selection and color averaging.
@@
Our approach, named  Im2SurfTex, addresses the above limitations with the introduction of a novel, optimization-free module that can be easily integrated to existing texture generation approaches. The module is trained to combine color information from multiple viewpoints to textures through a cross-attention mechanism, where for each surface point, several candidate image neighborhoods across different views are examined and combined back in texture space depending on local surface geometry, as encoded in 3D positions, normals and geodesic distances within these patches. In this manner, the attention mechanism captures local context based on surface (geodesic) proximity rather than relying solely on 3D Euclidean proximity that might correlate surface region textures far from each other in geodesic sense. Our module yields coherent textures efficiently, without requiring any slow optimization procedures. Our experiments indicate significant improvements in generated texture quality, measured by different scores, including FID [Heusel:2017:GANs], KID [Bińkowski:2018d:Demystifying], CLIP [Schuhmann:2022:Laion5b] metrics, when compared to alternatives.
@@ 2 Related Work
Classic texture generation methods mostly focused on example-based approaches [Wei:2009:STARtexture], including region-growing techniques [Efros:1999:texture, Wei:2000:Fast] and strategies leveraging local coherence [Ashikhmin:2001:Synthesizing, Tong:2002:BTF] to maintain consistency across synthesized textures. Patch-based methods [Efros:2001:Quilting, Kwatra:2003:Graphcuttex] synthesized textures using patch patterns from reference images, while other techniques [Kwatra:2005:TextureOptim, Han:2006:FastOptim] progressively refined synthesized texture based on optimization procedures. Another significant research direction involved directly generating textures on 3D surfaces [Turk:2001:Surfaces, Zhang:2006:vector, Fisher:2007:Design] by exploiting vector fields defined over the surface to seamlessly map textures onto complex geometries. All these example-based approaches were unable to capture texture variability, generate diverse textures, or handle diverse shapes.
@@
Our method builds upon diffusion models [Sohl:2015:Deep, Ho:2020:Denoising, Zhang:2023:T2I], which have demonstrated superior performance compared to GANs [Goodfellow:2014:GAN, Zhu:2017:unpaired] in image generation tasks [Rombach:2022:High, Nichol:2021:glide, Saharia:2022:imagen, Ramesh:2022:dalle, Zhang:2023:ControlNet]. Closely related to our approach are text-guided generation models for 3D object synthesis, where text-to-image diffusion models are used for distilling 3D objects as neural radiance fields [Mildenhall:2020:nerf, Kerbl:2023:3Dgaussians] via Score Distillation Sampling [Poole:2022:Dreamfusion, Wang:2023:score]. Following DreamFusion [Poole:2022:Dreamfusion], several approaches have been proposed [Lin:2023:Magic3D, Metzer:2023:latentnerf, Chen:2024:text2gs, Wang:2024:Prolificdreamer, Tsalicoglou:2024:textmesh, Shi:2024:mvdream]. However, these methods do not specifically target the task of texture generation.
@@
Initial efforts in texture generation via text-to-image diffusion models, such as Text2Tex [Chen:2023:Text2tex] and TEXTure [Richardson:2023:Texture], employed depth-conditioned diffusion models [Rombach:2022:High, Zhang:2023:ControlNet] to iteratively inpaint and refine the textures of 3D objects. Both methods start with a preset viewpoint, generating texture updates for corresponding regions of the 3D object by back-projecting depth-guided views. In Text2Tex, a coarse texture is progressively created by iterating over multiple viewpoints and refining the texture map based on high- surface-coverage viewpoints. This refinement applies a denoising diffusion process of moderate strength to preserve the texture’s original appearance while enhancing details. Similarly, TEXTure divides the texture map into distinct regions labeled as keep, refine, or generate, enabling selective refinement or generation of textures. Despite these efforts to achieve global consistency, these methods employ ad hoc thresholds to define the different shape regions and hand-engineered strategies for backprojection, often leading to seams between texture regions synthesized from different viewpoints.
@@
To address these issues, other methods such as TexFusion [Cao:2023:Texfusion] leverage latent diffusion to interlace diffusion and back-projection steps, producing 3D-aware latent images that are subsequently decoded and merged into a texture map. Similarly, SyncMVD [Liu:2024:SyncMVD] employs a latent texture map where all views are encoded at each denoising step, further enhancing consistency in geometry and appearance. TexGen [Huo:2024:TexGen] introduced a multi-view sampling and resampling framework that updates a UV texture map iteratively during denoising, aiming to reduce view discrepancies. Still, these methods rely on ad hoc blending masks or heuristics for aggregating texture information from different views, such as mere averaging or using the most front-facing view information for each texel. In contrast, our approach learns to aggregate color information from multiple views based on both geometry and texture information, promoting the generation of more coherent and seamless surface texture maps.
@@
A more recent approach, Paint3D [Zeng:2024:Paint3D], achieves impressive texture generation results, by adopting a two-stage texture generation strategy. In the first stage, a coarse texture is created by backprojecting views to texture space via the heuristic of using the most front-facing view information for each texel, as in previous methods. The second stage involves a refinement and inpainting process that utilizes a diffusion model in UV texture space, conditioned on a UV position map encoding the 3D adjacency information of texels. While this method directly encodes 3D geometric information into the texture map, it can still result in misaligned textures due to its employed heuristic during its coarse stage. The subsequent texture refinement steps often fail to fix the artifacts of the coarse stage, as demonstrated in our experiments. Another related method, MatAtlas [Ceylan:2024:Matatlas], incorporates a three-step denoising process with sequential operations and line conditions to preserve geometry and style consistency. However, MatAtlas employs an averaging heuristic for blending the final texture from the generated views, leading to inconsistencies or overly smooth surfaces, as also demonstrated in our experiments. TEXGen [Yu:2024:TEXGen] takes a different approach by directly training a large-scale diffusion model in the UV texture space, and integrating convolution operations in UV space with 3D-aware attention layers in their denoising network to achieve high-resolution texture synthesis. However, their method still faces challenges in maintaining cross-view consistency since the generated textures are conditioned on single-view images which are merely backprojected to the UV space to derive the initial partial texture maps used in their diffusion.
@@
In a concurrent work, MVPaint [Cheng:2024:MVPaint] introduces a multi-stage texture generation framework. In the initial stage, a latent texture map is employed during multi-view projection to create a synchronized texture across multiple views, similar to SyncMVD [Liu:2024:SyncMVD]. This is followed by an inpainting stage, where uncovered texture regions are filled using a dense colored point cloud extracted from the generated texture map. Colors are propagated to empty texels in a spatially aware manner using inverse distance weighting and normal similarity between neighboring points. Finally, a refinement stage upscales the texture map and smooths out seams through weighted color averaging among k-nearest neighbors in 3D space. Still, MVPaint relies on an averaging scheme to aggregate view information into texture space. In contract, our approach learns this aggregation by encoding both geometric and appearance information from multiple views to produce textures with greater consistency.
@@
Expanding beyond single object texture generation, InstanceTex [Yang:2024:Instancetex] focuses on texture generation for 3D scenes, employing a local synchronized multi-view diffusion strategy to improve local texture consistency across multiple objects. 3D Paintbrush [Cecatur:2024:Paintbrush] specializes in localized stylization of single objects, using cascaded score distillation to refine textures within specific object regions. These approaches differ from our method in scope: InstanceTex is tailored for stylistic consistency in large environments, while 3D Paintbrush targets localized edits.
@@
Figure 2: (Top) The Im2SurfTex, pipeline utilizes depth images and a text prompt to generate a number of candidate views (RGB images) for a given shape. The views are aggregated through a learned backprojection module that incorporates geometric information, such as 3D location, normals, angles between normals, and view vectors, as well as geodesic neighborhood information (bottom right) of shape points corresponding to pixels of the generated RGB images. The backprojection module integrates several cross-attention blocks (bottom left) used to infer texel features and colors from the appearance and geometric information gathered from relevant, non-background pixels across all available views. As some texels may remain uncolored, an inpainting and high-definition (HD) module is applied to refine the texture map following Paint3D [Zeng:2024:Paint3D].
@@ 3.1 Depth/edge map rendering & viewpoint selection
As done in several recent texture generation approaches [Zeng:2024:Paint3D, Liu:2024:SyncMVD, Ceylan:2024:Matatlas, Cheng:2024:MVPaint], the first step in our pipeline is to render the mesh into a set of depth maps $\{\mathbf{D}_{c}\}_{c=1}^{C}$ from various viewpoints, where $C$ is the total number of viewpoints. These maps are used as conditioning to guide the diffusion process to generate images consistent with the depth cues. There have been various strategies for viewpoint selection and diffusion model conditioning – in our paper, we experimented with two backbones: one based on Paint3D [Zeng:2024:Paint3D], and another based on MatAtlas [Ceylan:2024:Matatlas], briefly described below.
@@
Paint3D follows an iterative strategy of viewpoint selection, image generation, and backprojection of the generated images to texture space. First, a couple of $1024\times 512$ depth images are generated from the frontal and rear views of the shape and are concatenated in a $2\times 1$ grid. The grid is passed as input to a diffusion process that generates a corresponding $2\times 1$ grid of RGB images. The use of both views as input to the diffusion model helps with the view consistency [Zeng:2024:Paint3D]. The generated images are backprojected to the shape texture through a simple inverse UV mapping strategy – we discuss backprojection strategies, including ours, in Section 3.3. The next iteration proceeds with two side-wise viewpoints, from which both depth images and partially colored RGB images are rendered from the partially textured mesh. These are provided as a grid to another diffusion process, whose generated images are again backprojected to UV space. The process repeats for one more step where two other top- and bottom-wise viewpoints are used. In total, three iterations (total $C=6$ views), with two viewpoints processed at a time, yielded the best results in Paint3D. Our experiments with this backbone follow the same iterative procedure and viewpoints – we only modify the backprojection.
@@
MatAtlas [Ceylan:2024:Matatlas] follows a different viewpoint selection, diffusion conditioning, and view generation strategy. Initially, a set $400\times 400$ depth maps are rendered from viewpoints uniformly sampled from the viewing sphere, and are arranged in a $4\times 4$ grid. In addition, $16$ edge maps are created using the shape’s occluding and suggestive contours and are placed also in a grid. These two grids are used as input to the diffusion process that generates a $4\times 4$ grid of RGB images. These are backprojected and blended into the shape’s texture space (discussed in Section 3.3). The resulting partially textured shape is rendered from the same viewpoints, and the rendered RGB images along with added partial noise, are passed to a second diffusion process yielding an updated set of RGB views. These are backprojected to the shape’s texture space, yielding a sharper texture [Ceylan:2024:Matatlas]. In a third step, additional viewpoints are selected accessing shape regions not textured yet. The textured shape is rendered from these viewpoints and the rendered images are arranged in a grid processed through another diffusion process, which generates another set of RGB images. These are again backprojected to the final shape’s texture. In our implementation, we use $6$ initial viewpoints to render depth maps at resolution $512\times 512$, arranged in a $3\times 2$ grid (we do not make use of edge maps). The used viewpoints are the same as the ones used in Paint3D for more fair comparisons across the two backbones, and also because we observed that texture details are better preserved from the higher resolution depth maps. We replace the MatAtlas backprojection with ours – the rest of the pipeline follows MatAtlas.